THE ℓ^{2}-COHOMOLOGY OF ARTIN GROUPS

M. W. DAVIS AND I. J. LEARY

Abstract. For each Artin group we compute the reduced ℓ^{2}- cohomology of (the universal cover of) its “Salvetti complex”. This is a CW-complex which is conjectured to be a model for the classifying space of the Artin group. In the many cases when this conjecture is known to hold our calculation describes the reduced ℓ^{2}-cohomology of the Artin group.

1. Introduction

Suppose that G is a countable discrete group acting properly and cellularly on a CW-complex E with compact quotient. As in [14] or [16] one can define the reduced ℓ^{2}-cohomology group H^{i}(E) for each non-negative integer i. (We will give the definition in Section 3.) For each i, H^{i}(E) is a Hilbert space with an orthogonal G-action. When G is infinite, these Hilbert spaces tend to be infinite-dimensional unless they are zero. However, each H^{i}(E) is a “Hilbert G-module”, and to any Hilbert G-module V one may associate a real number, called its von Neumann dimension and denoted by dim_{G}(V). One can then define the ℓ^{2}-Betti numbers of E with respect to G by

ℓ^{2}b_{i}(E; G) := dim_{G} H^{i}(E).

If E is contractible and G acts freely on E (so that E/G is a classifying space for G), these ℓ^{2}-cohomology groups (and their von Neumann dimensions) are independent of the choice of E and so are invariants of G. In this case we will use the notation H^{i}(G) and ℓ^{2}b_{i}(G) to stand for H^{i}(E) and ℓ^{2}b_{i}(E; G) respectively.

ℓ^{2}-cohomology groups have proved difficult to compute directly. To date, most of the calculations have been based on two principles: (i) certain vanishing theorems, and (ii) Atiyah’s formula. A typical vanishing theorem will assert that for a certain class of groups G, H^{i}(G) is zero for i within a certain range. Atiyah’s formula states that in the case when G acts freely on E, the alternating sum of the ℓ^{2}-Betti numbers of G is equal to the ordinary Euler characteristic of E/G. In cases where H^{∗}(E) is known to vanish except in one dimension, this formula provides an exact calculation of the unique non-zero ℓ^{2}-Betti number. In this paper we compute the ℓ^{2}-Betti numbers ℓ^{2}b_{i}(G) when G is an Artin group. Surprisingly, there is a completely clean answer which is non-trivial in the sense that many of the ℓ^{2}b_{i}(G) are non-zero.

Date: February 22, 2002.

2000 Mathematics Subject Classification. Primary: 57M07; Secondary: 20F36, 20J05.

The first author was partially supported by NSF grant DMS-0104026. The second author acknowledges support from EPSRC (via grant GR/R07813), and from the Ohio State Mathematical Research Institute.

Associated to any Coxeter matrix (m_{ij}) on a finite set I, there is a Coxeter group W, an Artin group A and a simplicial complex L with vertex set I, called the “nerve” of the Coxeter matrix. (The definitions will be given in Section 5.) Coxeter groups are intimately connected to groups generated by reflections. For example, if n = Card(I), there is a standard representation of W as a linear reflection group on ℝ^{n}, so that there is a certain convex cone in ℝ^{n} with W acting properly on its interior Ω.

Artin groups have a related geometric interpretation. Consider the complexification of the W -representation described above. Then ℝ^{n} + iΩ is a a W -stable open convex subset of ℂ^{n} on which W acts properly. Moreover, W acts freely on the hyperplane complement

Y = (ℝ^{n} + iΩ) − ⋃_{r} H_{r},

where r ranges over all reflections in W and H_{r} denotes the intersection of the complex hyperplane fixed by r and (ℝ^{n} + iΩ). It is proved in [7] that there is a finite CW-complex X, called the “Salvetti complex” which is homotopy equivalent to Y/W. The complex X has one vertex and, for each k > 0, one k-cell for each (k − 1)-simplex of L. The 2-skeleton of X is the presentation complex associated to the standard presentation of A. It follows that π_{1}(Y/W) = π_{1}(X) = A. It is conjectured that the universal cover X̃ of X is contractible, or equivalently that X is a classifying space BA for A. Deligne [12] proved this conjecture in the case when W is finite. The conjecture is also known to be true when W is “right-angled”, when dim L ≤ 1, and in many other cases (see [6, 7]). Our actual calculation is of H^{i}(X̃). So our title is somewhat misleading: we only have a computation of the reduced ℓ^{2}-cohomology of A modulo the conjecture that X = BA. Theorem 1. For any Artin group A and any i, there is an isomorphism of Hilbert A-modules:

H^{i}(X̃) ≅ \overline{H}^{i−1}(L) ⊗ ℓ^{2}(A).

Corollary 2. ℓ^{2}b_{i}(X̃; A) = \overline{b}_{i−1}(L).

Here \overline{b}_{j}(L) denotes the ordinary “reduced Betti number” of L, i.e., \overline{b}_{j}(L) is the dimension of the reduced homology group \overline{H}_{j}(L; ℝ).

When W is finite, the Artin group A is said to be “spherical”. If A is spherical, then its centre always has an infinite cyclic subgroup (coming from the ℂ^{∗} -action on the hyperplane complement). For such groups the reduced ℓ^{2}-cohomology is known to vanish (see Section 4). Hence, in the spherical case, Theorem 1 is well-known. This vanishing result in the spherical case is the key ingredient for our proof in the general case.

In fact, a stronger result than Theorem 1 is true. In the following statement, the “standard homomorphism” from an Artin group to ℤ is the homomorphism that sends each Artin generator to 1 ∈ ℤ. Theorem 3. Suppose that A′ is a normal subgroup of the Artin group A that is contained in the kernel of the standard homomorphism A → ℤ. Let X′ → X be the corresponding covering space of X and let G = A/A′. Then H^{i}(X′) and \overline{H}^{i−1}(L) ⊗ ℓ^{2}(G) are isomorphic as Hilbert G-modules. In particular, ℓ^{2}b_{i}(X′; G) = \overline{b}_{i−1}(L).

Theorem 3 was suggested by a result of [15]. Acknowledgement. The authors wish to thank Tadeusz Januszkiewicz, Wolfgang Lück and Thomas Schick for numerous helpful conversations concerning the work in this paper.

2. Hilbert G-modules

Let G be a countable discrete group. The Hilbert space of squaresummable real-valued functions on G is denoted by ℓ^{2}(G). There is an orthogonal (right) action of G on ℓ^{2}(G), and a diagonal action on any direct sum of copies of ℓ^{2}(G). A Hilbert space V equipped with an orthogonal(right) G-action is called a (finitely generated) Hilbert Gmodule if it is G-equivariantly isomorphic to a G-stable closed subspace of the direct sum of finitely many copies of ℓ^{2}(G). A map f : V → V′ of Hilbert G-modules is by definition a G-equivariant bounded linear operator. The kernel of f : V → V′, denoted Ker(f), is a closed subspace of V, but the image, Im(f), need not be closed. The map f is weakly surjective if \overline{Im(f)} = V; f is a weak isomorphism if it is both injective and weakly surjective.

To each Hilbert G-module V it is possible to associate a non-negative real number dim_{G}(V), called its von Neumann dimension. The three most important properties of dim_{G}(−) are:

(1) dim_{G}(V) = 0 if and only if V = 0; (2) dim_{G}(V ⊕ V′) = dim_{G}(V) + dim_{G}(V′);

(3) dim_{G}(ℓ^{2}(G)) = 1.

It follows from (2) that for any map f : V → V′,

dim_{G}(V) = dim_{G}(Ker(f)) + dim_{G}(\overline{Im(f)}).

Using property (1), one can then deduce the following. Lemma 4. Suppose that V and V′ are Hilbert G-modules and that dim_{G}(V) = dim_{G}(V′). The following are equivalent:

(a) f is injective;

(b) f is weakly surjective;

(c) f is a weak isomorphism.

Proof. By definition, (c) is equivalent to the conjunction of (a) and (b). Thus it suffices to show that (a) and (b) are equivalent. But we have that dim_{G}(\overline{Im(f)}) = dim_{G}(V)−dim_{G}(Ker(f)), and so dim_{G}(Ker(f)) = 0 if and only if dim_{G}(\overline{Im(f)}) = dim_{G}(V) = dim_{G}(V′). □

Next suppose that C^{∗} = {(C^{n}, δ_{n})} is a cochain complex of Hilbert G-modules, i.e., the coboundary maps δ_{i} : C^{i} → C^{i+1} are G-equivariant bounded linear operators. The cohomology group

H^{i}(C^{∗}) := Ker(δ_{i})/ Im(δ_{i−1})

need not be a Hilbert space since Im(δ_{i−1}) is not necessarily closed in Ker(δ_{i}). However, the reduced cohomology groups defined as

H^{i}(C^{∗}) := Ker(δ_{i})/\overline{Im(δ_{i−1})}

are Hilbert G-modules. Lemma 5. Suppose that (C^{∗}, δ_{∗}) and (C′^{∗}, δ′_{∗}) are cochain complexes of Hilbert G-modules and that f : C^{∗} → C′^{∗} is a weak isomorphism of cochain complexes (i.e., f is a cochain map and for each i, f_{i} : C^{i} → C′^{i} is a weak isomorphism). Then the induced map H(f) : H^{i}(C^{∗}) → H^{i}(C′^{∗}) is also a weak isomorphism. In particular, H^{i}(C^{∗}) and H^{i}(C′^{∗}) are isometric Hilbert G-modules.

Proof. Let Z^{i} and B^{i} (resp. Z′^{i} and B′^{i}) denote the cocycles and coboundaries in C^{∗} (resp. C′^{∗}). Since f is a cochain map, f(Z^{i}) ⊆ Z′^{i} and f(B^{i}) ⊆ B′^{i}. Moreover, since f is continuous, f(\overline{B^{i}}) ⊆ \overline{B′^{i}}. Since f is injective this gives

dim_{G}(Z′^{i}) ≥ dim_{G}(Z^{i}) and dim_{G}(B′^{i}) ≥ dim_{G}(B^{i}). (1)

From the short exact sequences

0 → Z^{i} → C^{i} → B^{i+1} → 0
0 → Z′^{i} → C′^{i} → B′^{i+1} → 0

one obtains

dim_{G}(C^{i}) = dim_{G}(Z^{i}) + dim_{G}(B^{i+1}), and
dim_{G}(C′^{i}) = dim_{G}(Z′^{i}) + dim_{G}(B′^{i+1}) (2).

Since f is a weak isomorphism, one has that dim_{G}(C′^{i}) = dim_{G}(C^{i}). Hence

dim_{G}(C^{i}) = dim_{G}(C′^{i}) = dim_{G}(Z′^{i}) + dim_{G}(B′^{i+1})
≥ dim_{G}(Z^{i}) + dim_{G}(B^{i+1}) = dim_{G}(C^{i}),

which together with (1) implies that

dim_{G}(Z′^{i}) = dim_{G}(Z^{i}) and dim_{G}(B′^{i}) = dim_{G}(B^{i}). (3)

It follows from Lemma 4 that the maps f : Z^{i} → Z′^{i} and f : B^{i} → B′^{i} are weak isomorphisms. This implies that H(f) : H^{i}(C^{∗}) → H^{i}(C′^{∗}) is weakly surjective.

To show that H(f) is injective, let H be the orthogonal complement to \overline{B^{i}} in Z^{i}. Note that H is closed and G-stable. Now

dim_{G}(Z′^{i}) = dim_{G}(\overline{f (H + \overline{B^{i}})})
= dim_{G} (\overline{f(H)}) + dim_{G} (\overline{B′^{i}}) − dim_{G} (\overline{f(H)} ∩ \overline{B′^{i}})
= dim_{G}(Z′^{i}) − dim_{G} (\overline{f(H)} ∩ \overline{B′^{i}}) ,

where the third equality follows from (3). Hence dim_{G} (\overline{f(H)} ∩ \overline{B′^{i}}) =

0, and therefore \overline{f(H)} is a complementary subspace for \overline{B′^{i}}. Since f is injective when restricted to H, it follows that so is H(f). (This follows from the fact that any two weakly isomorphic Hilbert G-modules are in fact equivariantly isometric [14].) □

3. Cohomology with local coefficients and ℓ^{2}-cohomology

Suppose that Y is a connected CW-complex with fundamental group π and universal cover Ỹ. Suppose also that Y is of finite type (i.e., has finite skeleta.) Let C_{∗}(Ỹ) denote the chain complex of Ỹ, which is a chain complex of finitely generated free (left) ℤπ-modules. By a local coefficient system on Y we shall mean a (left) ℤπ-module. The cochain complex for Y with coefficients M is defined by

C^{∗}(Y; M) := Hom_{ℤπ}(C_{∗}(Ỹ), M).

Let Y′ → Y be a regular covering space corresponding to a normal subgroup π′ of π, and let G denote the quotient group π/π′. Then ℓ^{2}(G) may be viewed as a π-G-bimodule, where the left action of π is via the action of G. The ℓ^{2}-cochains on Y′ are defined by

ℓ^{2}C^{∗}(Y′) := Hom_{ℤG}(C_{∗}(Y′), ℓ^{2}(G))
= Hom_{ℤπ}(C_{∗}(Ỹ), ℓ^{2}(G))
= C^{∗}(Y; ℓ^{2}(G)).

Note that the orthogonal right action of G on ℓ^{2}(G) makes ℓ^{2}C^{∗}(Y′) a chain complex of Hilbert G-modules. The unreduced and reduced cohomology groups of ℓ^{2}C^{∗}(Y′) will be denoted ℓ^{2}H^{∗}(Y′) and H^{∗}(Y′) respectively. The ℓ^{2}-Betti numbers of Y′ with respect to G are then defined by ℓ^{2}b_{i}(Y′; G) := dim_{G} H^{i}(Y′).

In the case when Y′ is acyclic, C_{∗}(Y′) is a free resolution for ℤ over ℤG. Any two such resolutions are chain homotopy equivalent. Similarly, if Y′ is only rationally acyclic, the rational chain complex C_{∗}(Y′, ℚ) is a free resolution for ℚ over ℚG, and is unique up to chain homotopy equivalence. It follows that in the case when Y′ is rationally acyclic, the ℓ^{2}-cohomology groups and ℓ^{2}-Betti numbers depend only on G (and not on the particular choice of Y′). In this case we shall use the notation

ℓ^{2}H^{i}(G) := ℓ^{2}H^{i}(Y′),
H^{i}(G) := H^{i}(Y′),
ℓ^{2}b_{i}(G) := dim_{G} H^{i}(Y′).

(Of course, we have in mind the case when G = π and Y′ = Ỹ is contractible.)

4. A vanishing theorem

Let G be a group and Z a central subgroup of G. Any (left) Gmodule M may be viewed as a G-Z-bimodule where the right action of z ∈ Z is defined by mz = zm. This induces a right action of Z on H^{∗}(G; M). However one has: Proposition 6. With notation as above, the action of Z on H^{∗}(G; M) is trivial.

Proof. Let P_{∗} be a free resolution for ℤ over ℤG, so that H^{∗}(G; M) may be computed using the cochain complex Hom_{G}(P_{∗}, M). In terms of this cochain complex the right action of z ∈ Z on H^{∗}(G; M) sends f ∈ Hom_{G}(P_{i}, M) to zf, defined by zf(p) = z(f(p)) for all p ∈ P_{i}. It suffices to show that this map is chain-homotopic to the identity map. Since z is central, the map P_{∗} → P_{∗} given by p ↦ zp is a chain map, and this chain map lifts the identity map from ℤ to ℤ, so is chainhomotopic to the identity map of P_{∗}. But for all f ∈ Hom_{G}(P_{∗}, M) and all p ∈ P_{∗}, zf(p) = f(zp), and so the two actions of Z (on M and on P_{∗}) induce the same action on Hom_{G}(P_{∗}, M). □

Corollary 7. Let π be a group admitting a finite classifying space Y, let π′ be a normal subgroup of π, let G = π/π′, and suppose that the centre of π contains an element whose image in G has infinite order. Then H^{∗}(Y′) = {0}, where Y′ denotes the regular cover of Y corresponding π′.

Proof. Let Z denote an infinite cyclic subgroup of the centre of π with Z ∩π′ = {1}. The usual right action of Z on ℓ^{2}(G) and the action used in Proposition 6 agree, and hence the right action of Z on ℓ^{2}H^{∗}(Y′) is trivial. It follows that the action of Z on H^{∗}(Y′) is trivial. Since the image of Z in G is infinite, no non-zero element of ℓ^{2}(G) is fixed by the right action of Z. However, each H^{i}(Y′) is a Hilbert G-module, and so one obtains a contradiction unless each H^{i}(Y′) = {0}. □

Corollary 8. Suppose that G has a finite classifying space and that the centre of G contains an element of infinite order. Then the ℓ^{2} cohomology of G vanishes.

Proof. This is just the case G = π of Corollary 7. □

Other more general vanishing theorems for ℓ^{2}-cohomology may be found in [8, 16, 19, 20].

5. Coxeter groups and Artin groups

Let I be a finite set. A Coxeter matrix M = (m_{ij}) on I is an I-by-I symmetric matrix with entries in ℕ ∪ {∞} such that each m_{ii} = 1 and such that whenever i ≠ j, m_{ij} ≥ 2. Associated to M there is a Coxeter group denoted by W_{I} or W with generating set {s_{i}: i ∈ I}, with relations

(s_{i}s_{j})^{m_{ij}} = 1 for all (i, j) ∈ I × I.

There is also an Artin group A_{I} (or A) with generators {a_{i}: i ∈ I} and with a presentation given by the relations

a_{i}a_{j} ⋯ = a_{j}a_{i} ⋯

for all i ≠ j, where there are m_{ij} terms on each side of the equation. In both the Coxeter and Artin cases, the relation is interpreted as being vacuous if m_{ij} = ∞.

Since each of the Artin relators contains the same number of copies of a generator on each side, there is a homomorphism from any Artin group to the infinite cyclic group ℤ, defined by sending each Artin generator to 1. Call this the standard homomorphism from A to ℤ.

Let q : A → W denote the natural homomorphism sending a_{i} to s_{i}. There is a set-theoretic section for q denoted w ↦ a_{w} which may be defined as follows: if s_{i_{1}} ⋯ s_{i_{n}} is any word of minimal length for w in terms of the s_{i}, then a_{w} := a_{i_{1}} ⋯ a_{i_{n}}. As explained in [6, p. 602], it follows from Tits’ solution to the word problem for Coxeter groups that w ↦ a_{w} is well-defined.

Given a subset J of I, let M_{J} denote the minor of M whose rows and columns are indexed by J, and let W_{J} (resp. A_{J}) be the corresponding Coxeter group (resp. Artin group). It is known (cf. [2]) that the natural map W_{J} → W_{I} (resp. A_{J} → A_{I}) is injective, and hence W_{J} (resp. A_{J}) can be identified with the subgroup of W_{I} (resp. A_{I}) generated by {s_{i}: i ∈ J} (resp. {a_{i}: i ∈ J}).

Say that the subset J ⊆ I is spherical if W_{J} is finite. If this is the case, then the groups W_{J} and A_{J} are both called spherical. (Note that this differs from the terminology of [6], where spherical Artin groups were said to be “of finite type”.) The poset of spherical subsets of I is denoted by S(I) or simply S. The subposet of non-empty spherical subsets is an abstract simplicial complex which will be denoted L, and called the nerve of M (or the nerve of (W, {s_{i}: i ∈ I})). Thus the vertex set of L is I and a subset σ of I spans a simplex of L if and only if σ ∈ S if and only if W_{σ} is finite. (Greek letters such as σ or τ will be used to denote spherical subsets of I when viewed as simplices of L.)

Example 1. If W is finite, then L is a simplex of dimension Card(I)−1.

Suppose now that m_{ij} = ∞ for all i ≠ j, and let n = Card(I). In this case W is the free product of n copies of ℤ/2, A is a free group of rank n, and L is a 0-dimensional complex consisting of n points.

A Coxeter matrix is right-angled if each off-diagonal entry is either 2 or ∞. The associated Coxeter group and Artin group are also said to be right-angled.

For any subset J of I and w ∈ W_{I}, the following two conditions are equivalent:

(1) w is the (unique) element of shortest word length in wW_{J} ∈ W_{I}/W_{J}; (2) for each j ∈ J, l(ws_{i}) = l(w) + 1 (where l(w) denotes the word

length of w). If w satisfies either condition, then it is called J-reduced. Moreover, if w is J-reduced and v ∈ W_{J}, then l(wv) = l(w) + l(v) (cf. [2, Ex. 3, p. 37]). For each pair J, J′ of subsets of I with J ⊆ J′, let W_{J′}^{J} denote the set of J-reduced elements of W_{J′}.

We recall now a formula from [22, 11]. Given a pair of spherical subsets τ, σ with τ ≤ σ, define an element T_{σ}^{τ} ∈ ℤA_{σ} by the formula

T_{σ}^{τ} := ∑_{w∈W_{σ}^{τ}}(−1)^{l(w)}a_{w}.

Lemma 9. For any three spherical subsets, τ, σ, ρ ∈ S with τ ≤ σ ≤ ρ, we have

T_{ρ}^{σ} · T_{σ}^{τ} = T_{ρ}^{τ}.

Proof. If u ∈ W_{ρ}^{σ} and v ∈ W_{σ}^{τ}, then l(uv) = l(u) + l(v) It follows that a_{u}a_{v} equals a_{uv}. Since, for each i ∈ τ, l(uvs_{i}) = l(u) + l(vs_{i}) = l(u) + l(v) + 1 = l(uv) + 1, the element uv is τ-reduced. Hence, as u ranges over W_{ρ}^{σ} and v ranges over W_{σ}^{τ}, uv ranges over W_{ρ}^{τ}. Therefore,

T_{ρ}^{σ} · T_{σ}^{τ} = ( ∑_{u∈W_{ρ}^{σ}}(−1)^{l(u)}a_{u}) ( ∑_{v∈W_{σ}^{τ}}(−1)^{l(v)}a_{v})

= ∑_{w∈W_{ρ}^{τ}}(−1)^{l(w)}a_{w} = T_{ρ}^{τ}.
□

6. Some homological algebra

Let L be a finite abstract simplicial complex. Later we shall assume that L is ordered, in the sense that there is a given linear ordering of its vertex set. The face category F(L) of L is defined to be the category whose objects are the simplices of L, with one morphism from τ to σ whenever τ is a face of σ. The augmented face category F^{+}(L) has the same objects and morphisms as F(L), but also has one extra object (the “−1-simplex”, denoted ∅) equipped with one morphism to each object. If L has vertex set I, each of F(L) and F^{+}(L) is isomorphic to a full subcategory of the category of subsets of I. A cosheaf on L with values in a category C is a covariant functor F from F^{+}(L) to C. In the cases of interest to us, C will be either abelian groups or Hilbert G-modules. A homomorphism of cosheaves on L is given by ϕ : F → F′, a natural transformation from F to F′.

If τ ≤ σ are faces of L (possibly including the −1-simplex), let ι_{σ}^{τ} denote the morphism in F(L) from τ to σ. The morphisms of F(L) are generated by those ι_{σ}^{τ} such that τ is a codimension one face of σ, and all relations between the morphisms are consequences of the relations

ι_{τ′}^{σ′}ι_{σ}^{τ′} = ι_{τ}^{σ′}ι_{σ}^{τ}, (4)

where σ′ is a codimension two face of σ, and τ, τ′ are the two codimension one faces of σ containing σ′. To define a cosheaf F on L it suffices to specify the objects F (σ), and morphisms f_{σ}^{τ} = F (ι_{σ}^{τ}) : F (σ) → F (σ) for all codimension one pairs, so that relations (4) are satisfied for all codimension two pairs.

Now suppose that L is ordered. Then for any n ≥ 0 the vertices of an n-simplex form an ordered set isomorphic to 0 < 1 < ⋯ < n. For any 0 ≤ i ≤ n and any n-simplex σ, the ith face of σ is defined to be the (n − 1)-simplex spanned by all vertices of σ except the ith. If one writes ∂_{i} = ι_{σ}^{τ}, where τ is the ith face of σ, then the relations between the morphisms become the familiar “cosimplicial identities” as in [24] 8.1.

A cosheaf F of abelian groups on an ordered simplicial complex L gives rise to a cochain complex C^{∗}(L; F) defined as follows: C^{n} = 0 for n < −1, and for n ≥ −1,

C^{n} = ⨁_{τ∈L^{n}} F (τ),

where the indexing set is the set of n-simplices of L. Under the natural isomorphism

Hom(C^{n}, C^{n+1}) ≅ ⨁_{τ∈L^{n}, σ∈L^{n+1}} Hom(F (τ), F (σ)),

the coboundary map d : C^{n} → C^{n+1} corresponds to the matrix (d_{τσ}), where d_{τσ} = 0 unless τ is a face of σ, and is equal to (−1)^{i}F (ι_{σ}^{τ}) if τ is the ith face of σ. Example 2. For any abelian group N, and any simplicial complex L, there is a constant cosheaf on L denoted by \underline{N}, which sends each object of F^{+}(L) to N and each morphism to the identity map. The cohomology of the corresponding cochain complex is isomorphic to the reduced cohomology of L with coefficients in N:

H^{∗}C^{∗}(L; \underline{N}) = \overline{H}^{∗}(L; N).

Now suppose that M is a Coxeter matrix with nerve L and Artin group A, and let N be a ℤA-module. Define a cosheaf N on L by sending each object to N, and for each τ < σ sending the morphism ι_{σ}^{τ} to left multiplication by T_{σ}^{τ}. (Lemma 9 implies that this does define a cosheaf.)

One of the key observations of this paper is: Lemma 10. For any Coxeter matrix, M, with L, A and N as above, there is a cosheaf homomorphism ϕ : \underline{N} → N defined by

ϕ(σ) : n ↦ T_{σ}^{∅} · n.

Proof. Immediate from Lemma 9. □

Corollary 11. Let N be a ℤA-module upon which T_{σ}^{∅} acts isomorphically for each σ ∈ L. Then

H^{∗}C^{∗}(L; N) ≅ \overline{H}^{∗}(L; N)

Proof. In this case the map described in Lemma 10 is an isomorphism of cosheaves, and so induces an isomorphism of cochain complexes C^{∗}(L; \underline{N}) → C^{∗}(L; N). □

7. The Salvetti Complex

Let M be a spherical Coxeter matrix on a set I of cardinality n. By definition the associated Coxeter group W is finite, and so admits a standard representation as an orthogonal reflection group on ℝ^{n}. In this case, Salvetti [22, 23] defined a regular CW-complex X′ which is equivariantly homotopy equivalent to Y, the complexification of ℝ^{n} minus the hyperplanes fixed by reflections in W. Each closed cell of X′ is naturally identified with a face of a certain convex polytope associated to W (namely the convex hull of a generic W -orbit in ℝ^{n}). Such a polytope is called a “Coxeter cell”. The group W acts freely on X′. The quotient space, denoted by X_{I} or X, is called the Salvetti complex, and has fundamental group A.

As mentioned in the introduction, the definition of X was extended to arbitrary Coxeter matrices in [7]. As before, there is a regular W - CW-complex X′ such that each closed cell is a Coxeter cell and such that X′ is homotopy equivalent to a suitably defined hyperplane complement Y. The quotient space X = X′/W is the Salvetti complex, and has fundamental group A. The space X has one open cell e_{σ} for each spherical subset σ of I. Moreover the closure of e_{σ} in X is the subcomplex X_{σ} corresponding to the spherical Artin group A_{σ}. Example 3. Let 𝕋^{n} denote the usual cubical structure on the n-torus, with one k-dimensional cell for each subset of I = {1, …, n} of cardinality k. In the case when W = (ℤ/2)^{n}, then A = ℤ^{n}, and X = 𝕋^{n}.

More generally, if A is right-angled with generating set I of cardinality n, then X is a subcomplex of 𝕋^{n}, consisting of those cells of 𝕋^{n} that correspond to spherical subsets of I.

It is proved in [7] that the poset of cells in X′ can be identified with W × S, where the partial order given by incidence of faces is given by (w, τ) < (v, σ) if and only if τ < σ and v^{−1}w ∈ W_{σ}^{τ}. Henceforth we identify a cell with the corresponding element of W × S. The cell corresponding to (v, σ) is a Coxeter cell associated to W_{σ}; its vertex set is the set {(vu, ∅): u ∈ W_{σ}^{∅}} and its dimension is Card(σ).

An edge or 1-cell of X′ has the form (w, {i}) for some w ∈ W and i ∈ I. Orient this edge by declaring (w, ∅) to be its initial vertex and (ws_{i}, ∅) to be its terminal vertex. Since the W -action preserves the edge orientations, this gives rise to an orientation of the edges of X = X′/W. Let {e_{i}: i ∈ I} denote the set of edges of X with their induced orientation. Since X has only one vertex, each e_{i} gives rise to an oriented loop. By definition the generator a_{i} of the Artin group A = π_{1}(X) is the homotopy class of the oriented loop around e_{i}.

A vertex x of a cell C in X′ is called a top vertex (resp. bottom vertex) of C if each edge of C that contains x points away from x (resp. towards x). Each cell (v, σ) of X′ has a unique top vertex, (v, ∅), and a unique bottom vertex (vu_{σ}, ∅), where u_{σ} denotes the element of longest length in the finite Coxeter group W_{σ}.

Let X̃ denote the universal cover of X (or equivalently of X′). The orientations of the edges of X′ lift to orientations on the edges of X̃. Each cell of X̃ then also has a unique top vertex and a unique bottom vertex. The 1-skeleton of X̃ is an oriented graph and each edge is labelled by an element of {a_{i}: i ∈ I} so that if e is an oriented edge labelled by a_{i} whose initial vertex is x, then the terminal vertex of e is x · a_{i}.

Now fix a lift x̃ ∈ X̃ of the vertex (1, ∅) of X′. This gives an identification of the vertex set of X̃ with A and of the 1-skeleton of X̃ with the Cayley graph of the standard presentation of A.

For each σ ∈ S, let ẽ_{σ} be the lift of the cell (1, σ) whose top vertex is x̃. Using the A-action we get an identification of the poset of cells in X̃ with A × S. The partial order on A × S corresponding to “is a face of” is defined by (b, τ) < (a, σ) if and only if b = aa_{u} for some u ∈ W_{σ}^{τ}. Thus for a given cell (a, σ) of X̃, its set of faces of type τ is {(aa_{u}, τ): u ∈ W_{σ}^{τ}}.

Once we have fixed an ordering on I there is an induced orientation for each cell of X̃: a cell C with top vertex x is oriented so that the edges at x in their natural order give an oriented basis for the tangent space of C at x (recall that C may be viewed as a convex polytope in ℝ^{n} for some n). Suppose that C is of type σ, σ = {i_{1} < ⋯ < i_{k}} and that D is a codimension one face of type τ = σ − {i_{j}}. Suppose also that D has the same top vertex as C. In this case the natural orientation of D and its orientation induced by the orientation on C (by taking an outward pointing normal) differ by the sign ϵ(σ, τ) defined by ϵ(σ, τ) = (−1)^{j−1}.

If (v, σ) is a cell of X′, then the (right) action of W_{σ} on the cell (induced by the action on its vertex set) is such that the action of u ∈ W_{σ} changes the sign of the orientation by (−1)^{l(u)}. Hence, we have a formula for the boundary maps in C_{∗}(X̃):

∂(aẽ_{σ}) = ∂(a, σ) = ∑_{τ} ϵ(σ, τ) ∑_{u∈W_{σ}^{τ}}(−1)^{l(u)}(aa_{u}, τ)
= a ∑_{τ} ẽ_{τ}T_{σ}^{τ} · ϵ(τ, σ).

where the summation is over all codimension one faces τ of σ. Remark. Note the similarity with the formula occuring in Lemma 9. An alternative proof of Lemma 9 can be given by considering the oriented incidence relation between cells in X̃ of types τ, σ and ρ, where τ ⊆ σ ⊆ ρ.

The cells ẽ_{σ} for σ ∈ S form a ℤA-basis for the cellular chain groups of X̃. In matrix terms, the ith cellular chain group of X̃ may be viewed as consisting of row vectors over ℤA, with the standard basis elements corresponding to the cells ẽ_{σ}, where σ ranges over the subsets of S of size i. The ℤA-action is the standard left action, and the boundary map in the chain complex is given by right multiplication by a matrix whose entry in the (σ, τ)-position is ϵ(τ, σ)T_{σ}^{τ} if τ ⊆ σ and is zero otherwise.

Now suppose that N is any ℤA-module. After identifying C_{i}(X̃) with a space of row vectors with entries in ℤA as above, it is natural to identify Hom_{A}(C_{i}(X̃), N) with a space of column vectors (of the same length) with entries in N, so that the evaluation map corresponds to multiplying a row vector and a column vector to produce a single element of N. In these terms, the coboundary map in the cochain complex Hom_{A}(C_{∗}(X̃), N) is described as left multiplication by the same matrix over ℤA as used previously to describe the boundary map in C_{∗}(X̃). This gives the following theorem.

Theorem 12. For any A-module N, there is an isomorphism between C^{∗}(X; N) = Hom_{A}(C_{∗}(X̃), N) as defined at the beginning of Section 3 and C^{∗−1}(L; N) as defined in Section 6.

In the spherical case, a statement equivalent to the above theorem is contained in [23, 11], and it is used to compute some cohomology. Remark. There is a simple description of the CW-complex X. Firstly, for each σ ∈ S we describe a CW-complex X_{σ}. Start with a Coxeter cell C_{σ} of type W_{σ}. For each τ < σ identify C_{τ} with the face of C_{σ} of type τ which has the same top vertex as C_{σ}. Now for each τ < σ and each u ∈ W_{σ}^{τ} glue together C_{τ} and uC_{τ} via the homeomorphism induced by u. The result is X_{σ}. To construct X, start with the disjoint union of the X_{σ} for σ ∈ S, and then use the natural maps to identify X_{τ} with a subcomplex of X_{σ} whenever τ < σ. Remark. In section 3 of [6] it is shown how to associate a “simple complex of groups” A to a Coxeter matrix. The underlying poset is S. To each σ ∈ S one associates the Artin group A_{σ}, and for each τ < σ, the associated homomorphism A_{τ} → A_{σ} is the natural inclusion. There is a natural projection from (the barycentric subdivision of) X to the geometric realization of S such that the inverse image of the vertex corresponding to σ is a copy of X_{σ}. Since it follows from [12] that X_{σ} ≃ BA_{σ}, it follows that X is an aspherical realization of the complex of groups A in the sense of [17]. Thus X is homotopy equivalent to the classifying space BA of A (which is defined as a homotopy colimit of BA_{σ} over the category S). The main conjecture of [6] is that BA (or equivalently X) is aspherical. This has been proved in many cases: for example, it holds if the simplicial complex L is a flag complex (e.g., if the Coxeter matrix is either spherical or right-angled) and it holds whenever the dimension of L is at most 1.

8. Cohomology with generic coefficients

This section represents a slight digression, although the methods used are similar to those that will be applied to ℓ^{2}-cohomology in later sections.

Let A be an Artin group with Artin generators {a_{i}: i ∈ I}, let G be the abelianization of A, and let α_{i} be the image of a_{i} in G. In the case when m_{ij} is even, the abelianization of the Artin relator between a_{i} and a_{j} is trivial, and in the case when m_{ij} is odd the abelianization of the Artin relator is α_{i} = α_{j}. Hence one sees that G is free abelian, of rank equal to the number of components of the graph with vertex set I and an edge joining i to j if and only if m_{ij} is odd. In particular, G is never trivial. Let m denote the rank of G, and (after changing the ordering on I), suppose that α_{1}, …, α_{m} freely generate G.

Now suppose that k is a field of characteristic zero, and that N is a kA-module whose underlying k-vector space is 1-dimensional. The action of A on N factors through G, a free abelian group of rank m, and so N is classified up to isomorphism by an m-tuple

Λ = (λ_{1}, …, λ_{m}) ∈ (k^{∗})^{m},

where k^{∗} = k − {0} and for 1 ≤ i ≤ m, the generator α_{i} acts on N as multiplication by λ_{i}. As an illustration of the methods we shall apply to ℓ^{2}-cohomology, we shall compute the cohomology of the Salvetti complex X with local coefficients N for “generic N”, i.e., for N corresponding to elements of a dense Zariski open subset of (k^{∗})^{m}. Theorem 13. For a generic 1-dimensional kA-module N, there is an isomorphism

\overline{H}^{∗−1}(L; k) ≅ H^{∗}(X; N)

Proof. Each side of the supposed isomorphism is isomorphic to the cohomology of L with coefficients in a certain cosheaf. The left hand side is:

\overline{H}^{∗}(L; k) ≅ \overline{H}^{∗}(L; N) ≅ H^{∗}C^{∗}(L; \underline{N})

and by Theorem 12, the cochain complex for the right hand side is:

C^{∗}(X; N) ≅ C^{∗−1}(L; N).

It therefore suffices to show that for generic N, the cosheaf homomorphism ϕ : \underline{N} → N described in Lemma 10 is an isomorphism. Equivalently, it suffices to show that for each simplex σ of L, multiplication by T_{σ}^{∅} is an isomorphism for generic N.

For Λ = (λ_{1}, …, λ_{m}) ∈ (k^{∗})^{m}, let ψ′_{Λ} : ℤG → k be defined by ψ′_{Λ}(a_{i}) = λ_{i}, and let ψ_{Λ} : ℤA → k be the composite of the projection ℤA → ℤG and ψ′. For each σ ∈ L, the set of Λ such that ψ_{Λ}(T_{σ}^{∅}) = 0 is a closed subset of (k^{∗})^{m}. The union over all σ of these closed sets is not the whole of (k^{∗})^{m}, since in the case when Λ = (−1, −1, ⋯, −1), ψ_{Λ}(T_{σ}^{∅}) = Card(W_{σ}). Since (k^{∗})^{m} is connected, it follows that the set of Λ for which each ψ_{Λ}(T_{σ}^{∅}) ≠ 0 is open and dense in (k^{∗})^{m} as required. □

Remark. The theorem does not hold for arbitrary 1-dimensional coefficients, as can be seen by comparing the case of “generic” 1-dimensional coefficients N with the trivial 1-dimensional module k. As a first example, consider H^{1}. For any A, H^{1}(A; k) is naturally isomorphic to Hom(A, k), which has dimension equal to the rank of the abelianization of A. In particular it is non-zero for every Artin group, whereas for A of spherical type and generic N, we have seen that H^{∗}(A; N) = {0}. (Note that in the case when A is of spherical type, H^{∗}(A; k) is computed in [4].)

As a further example, consider the case when A is right-angled. In this case it may be shown that H^{∗}(A; k) is isomorphic to the “exterior face ring of L”, defined by taking the exterior algebra over k with generators {x_{i}: i ∈ I} with each x_{i} in degree 1, and adding in the relation x_{i_{1}} ⋯ x_{i_{m}} = 0 whenever i_{1}, …, i_{m} do not form the vertex set of a simplex of L [13, 18]. In particular, the rank of H^{i}(X; k) = H^{i}(A; k) is equal to the number of (i − 1)-simplices in L for each i > 0.

9. Reduced ℓ^{2}-cohomology

Each Artin group of spherical type has a non-trivial centre. If A is a spherical Artin group and w is the longest element of the corresponding Coxeter group, then the element (a_{w})^{2} is in the centre of A (c.f. [12]). Note that the expression for this element contains only positive powers of the Artin generators, and so in particular its image under the standard homomorphism A → ℤ is non-zero. Lemma 14. Suppose that A is an Artin group of spherical type, and that A′ is a normal subgroup of A contained in the kernel of the standard homomorphism. Let X′ → X be the corresponding covering space of X and let G = A/A′. For each i, the Hilbert G-module H^{i}(X′) is the zero module.

Proof. In view of the remarks preceding the statement of the lemma, this is a special case of Corollary 7. □

Proposition 15. Let H ≤ G, and let ξ ∈ ℤH. Then left multiplication by ξ induces a weak isomorphism of ℓ^{2}(H) if and only if it induces a weak isomorphism of ℓ^{2}(G).

Proof. Consider H, defined to be the set of all functions on G that are square-summable on each right coset Hg of H. As a left ℤHmodule, H is isomorphic to a direct product of copies of ℓ^{2}(H) indexed by the coset space H\G. One has that ℓ^{2}(H) ≤ ℓ^{2}(G) ≤ H, and so multiplication by ξ is injective as a self-map of ℓ^{2}(H) if and only if it is injective as a self-map of ℓ^{2}(G). The corresponding assertion with “weak isomorphism” in place of “injective” follows from Lemma 4. □

Lemma 16. Suppose that A = A_{σ} is an Artin group of spherical type, and that A′ is a normal subgroup of A contained in the kernel of the standard homomorphism A → ℤ. Let G = G_{σ} = A/A′, and let ϕ_{σ} : ℓ^{2}(G) → ℓ^{2}(G) denote left multiplication by the element T_{σ}^{∅} ∈ ℤA_{σ}. Then ϕ_{σ} is a weak isomorphism.

Proof. Let π denote the quotient map from A_{σ} to G_{σ}. For any τ < σ, let i_{τ} denote the inclusion of A_{τ} in A_{σ}, and define a subgroup G_{τ} ≤ G_{σ} by G_{τ} = π ∘i_{τ}(A_{τ}). The composite of i_{τ} and the standard homomorphism for A_{σ} is equal to the standard homomorphism for A_{τ}, and so the pair (A_{τ}, G_{τ}) satisfy the hypotheses of the Lemma.

The lemma is trivially satisfied if Card(σ) = 0, since T_{∅}^{∅} = 1. Hence by induction on Card(σ), we may assume that for any τ < σ, the map ϕ_{τ} : ℓ^{2}(G_{τ}) → ℓ^{2}(G_{τ}) is a weak isomorphism. Here ϕ_{τ} is defined to be left multiplication by the element T_{τ}^{∅} ∈ ℤA_{τ}. By Proposition 15, we may assume that ϕ_{τ} : ℓ^{2}(G_{σ}) → ℓ^{2}(G_{σ}) is also a weak isomorphism.

Now let L be the simplex with vertex set σ, where n = Card(σ). Let N denote ℓ^{2}(G_{σ}) = ℓ^{2}(G), and consider the cosheaves of Hilbert G-modules on L that were denoted by \underline{N} and N in Section 6. Note that in this case, C^{∗}(L; \underline{N}) and C^{∗}(L; N) are cochain complexes of Hilbert G-modules. Since the reduced cohomology of L is trivial, the cochain complex C^{∗}(L; \underline{N}) is exact. By Theorem 12 the cochain complex C^{∗}(L; N) (with a shift of degree) may be used to compute the cohomology of A with coefficients in ℓ^{2}(G), and so by Lemma 14, it follows that C^{∗}(L; N) is weakly exact.

Now consider the cosheaf homomorphism ϕ : \underline{N} → N defined in Lemma 10, and the induced map

ϕ^{∗} : C^{∗}(L; \underline{N}) → C^{∗}(L; N).

In degree i, each of C^{i}(L; \underline{N}) and C^{i}(L; N) is isomorphic to a direct sum of copies of ℓ^{2}(G) indexed by the subsets τ ≤ σ with Card(τ) = i. With respect to these bases, the map ϕ^{i} is given by a diagonal matrix, whose (τ, τ)-entry is the map ϕ_{τ}. By induction, this map is a weak isomorphism in each degree except possibly n, the top degree. Consider the following commutative diagram, consisting of the right-hand ends of the two cochain complexes.

⟶ C^{n−1}(L; \underline{N}) ⟶^{δ′} C^{n}(L; \underline{N})
↓^{ϕ_{n−1}} ↓^{ϕ_{n}}
⟶ C^{n−1}(L; N) ⟶^{δ} C^{n}(L; N)

This can be rewritten as follows, where τ runs over the subsets of σ of size n − 1.

⟶ ⨁_{τ} ℓ^{2}(G) ⟶^{δ′} ℓ^{2}(G)
↓^{⨁_{τ} ϕ_{τ}} ↓^{ϕ_{σ}}
⟶ ⨁_{τ} ℓ^{2}(G) ⟶^{δ} ℓ^{2}(G)

The coboundary map δ′ is surjective, the coboundary map δ is weakly surjective, and by induction the left-hand vertical map is a weak isomorphism. It follows that the right-hand vertical map, ϕ_{σ} : ℓ^{2}(G) → ℓ^{2}(G) is weakly surjective. The claim now follows from Lemma 4. □

Corollary 17. Let A be any Artin group, let A′ be a normal subgroup of A contained in the kernel of the standard homomorphism A → ℤ, let G = A/A′, and let N = ℓ^{2}(G). Then the map ϕ : \underline{N} → N defined in Lemma 10 is a weak isomorphism of cosheaves of Hilbert G-modules on the simplicial complex L.

Proof. On each object σ ∈ S, the map ϕ(σ) : \underline{N}(σ) → N (σ) is equal to left multiplication by T_{σ}^{∅}. By Lemma 16 and Proposition 15, each of these maps is a weak isomorphism. □

The next Corollary is immediate from the above.

Corollary 18. With notation as in Corollary 17, the induced map of cochain complexes

ϕ^{∗} : C^{∗}(L; \underline{N}) → C^{∗}(L; N)

is a weak isomorphism.

The proofs of Theorems 1 and 3 are now completed by combining the description of H^{(}X′) as the reduced cohomology of C^{∗}(L; N), given in Theorem 12, with Corollary 18 and Lemma 5.

10. Closing remarks

In this section we make some remarks concerning ℓ^{p}-cohomology for p ≠ 2, and describe an alternative method for proving Theorems 1 and 3, using a spectral sequence argument. In a subsequent paper with T. Januszkiewicz we plan to use this technique to calculate the ℓ^{2}-cohomology of the universal cover of the Salvetti complex associated to an arbitrary real hyperplane arrangement. To simplify the notation, we shall discuss this proof for Theorem 1 but not for Theorem 3.

10.1. A spectral sequence argument. Recall from Section 7 that for a general Artin group A, the Salvetti complex X may be expressed as a union of subcomplexes X_{σ}, where σ ranges over the spherical subsets of the Artin generating set, and X_{σ} is a copy of the Salvetti complex for A_{σ}. Let Y_{σ} be the subcomplex of X̃, the universal cover of X, consisting of lifts of cells of X_{σ}. Each Y_{σ} is a free A-CW-complex. By Corollary 8 and Proposition 15, the reduced ℓ^{2}-cohomology of Y_{σ} vanishes for σ ≠ ∅. In contrast, Y_{∅} is just the 0-skeleton of X̃, which consists of a single free G-orbit of cells. Hence the reduced ℓ^{2}-cohomology of Y_{∅} is a copy of ℓ^{2}(G) in degree zero.

One may construct a “Mayer-Vietoris double chain complex” C_{∗,∗} for X̃ expressed as the union of the Y_{σ}. Let L denote the set of maximal simplices in L, define Y (∅) = X̃, and for ∅ ≠ S ⊆ L, define Y (S) = ⋂_{σ∈S} Y_{σ}. Define C_{∗,0} to be the cellular chain complex for X̃, and for j > 0 define C_{∗,j} to be the direct sum over the subsets S ⊆ L of size j of the cellular chain complex for Y (S). The boundary maps in C_{∗,∗} of degree (−1, 0) are the boundary maps in the chain complexes C_{∗}(Y (S)), and the boundary maps of degree (0, −1) are given by the matrices whose (S, T)-entry is ϵ(S, T) times the map induced by the inclusion of Y (S) in Y (T), where ϵ(S, T) = (−1)^{i} if T is obtained from S by omitting the ith element of S (for some fixed ordering of L). By construction, this double complex has trivial homology, since the boundary map of degree (0, −1) is exact. Since each C_{i,j} is free, it follows that for each i, the chain complex C_{i,∗} is split exact.

Now suppose that N is a ℤA-module such that for each σ ∈ L, H^{∗}(A_{σ}; N) = {0}. Define a double cochain complex by

E_{0}^{∗,∗} = Hom_{A}(C_{∗,∗}, N),

and let E_{∗}^{∗,∗} denote the spectral sequence in which the differential d_{0} is induced by the boundary map of degree (−1, 0) on C_{∗,∗}. The boundary map on E_{0}^{∗,∗} of degree (0, 1) is exact, and so the homology of the total complex for E_{0}^{∗,∗} is zero. It follows that the E_{∞}-page of the spectral sequence is identically zero. The E_{1}-page has E_{1}^{i,0} ≅ H^{i}(X; N), E_{1}^{i,j} = {0} if both i > 0 and j > 0, and E_{1}^{0,j} is isomorphic to a direct sum of copies of N, indexed by those j-element subsets of L such that the intersection of the corresponding simplices of L is empty.

The structure of the E_{1}-page implies that E_{2}^{i,j} = E_{1}^{i,j} for i > 0, and it may be shown that E_{2}^{0,j} ≅ \overline{H}^{j}(L; N). To see this, note that the cochain complex E_{1}^{0,∗} embeds as a subcomplex in an exact complex C^{∗}, where C^{j} is isomorphic to a direct sum of copies of N indexed by all j-element subsets of L, and the quotient complex C^{∗}/E_{1}^{0,∗} is isomorphic to the augmented cochain complex for the nerve of the covering of L by the elements of L, shifted in degree by 1. The long exact sequence in cohomology coming from this short exact sequence of cochain complexes gives the claimed isomorphism.

Since the E_{∞}-page is identically zero, and the only non-zero groups on the E_{2}-page are those E_{2}^{i,j} for which either i = 0 or j = 0, it follows that the differential d_{r} must be an isomorphism from E_{r}^{0,r−1} = E_{2}^{0,r−1} to E_{r}^{r,0} = E_{1}^{r,0}. Thus we obtain Theorem 19. For any ℤA-module N such that H^{∗}(A_{σ}; N) = {0} for all non-trivial spherical subgroups A_{σ} ≤ A, there is an isomorphism

H^{∗}(X; N) ≅ \overline{H}^{∗−1}(L; N).

One corollary of this is a version of Theorem 13 Corollary 20. Let k be a field of characteristic zero, and let N be a 1-dimensional kA-module, such that for each non-trivial spherical subgroup A_{σ}, the centre of A_{σ} acts non-trivially on N. Then H^{∗}(X; N) ≅ \overline{H}^{∗−1}(L; N).

Proof. By Proposition 6, the action of the centre Z of A_{σ} on H^{∗}(A_{σ} : N) is trivial. But as a kZ-module H^{i}(A_{σ}; N) is isomorphic to a submodule of a direct sum of copies of N. The only possibility is that H^{∗}(A_{σ}; N) = {0}. Hence Theorem 19 may be applied. □

Unfortunately, Theorem 19 does not apply directly to the case when N = ℓ^{2}(A), since we have only that the reduced ℓ^{2}-cohomology groups H^{∗}(A_{σ}) are equal to {0}. One way around this difficulty is to use W. Lück’s equivalence of categories (for any discrete group G) between the category of Hilbert G-modules and the category of projective modules for the group von Neumann algebra N (G), as explained in chapter 6 of [20]. The von Neumann dimension is defined for projective N (G)-modules in such a way that it is preserved by the above equivalence of categories. Furthermore, von Neumann dimension extends to a dimension function on all (finitely generated) N (G)-modules so that it is additive on extensions, and such that dim_{G}(M) = 0 if and only if Hom_{G}(M, N (G)) = {0}. For any CW-complex Y with fundamental group G, the von Neumann dimension of the ordinary homology group H_{i}(Y; N (G)) is equal to the von Neumann dimension of H^{i}(Y). Hence the following statement is a weak version of Theorem 1. Theorem 21. For any Artin group A and any i, the von Neumann dimensions of H_{i}(X; N (A)) and \overline{H}_{i−1}(L; N (A)) ≅ N (A) ⊗ \overline{H}_{i−1}(L) are equal.

Proof. Use the homology version of the Mayer-Vietoris spectral sequence described above. In more detail, define a double chain complex

E_{∗,∗}^{0} = N (A) ⊗_{ℤA} C_{∗,∗}.

The corresponding spectral sequence has the following properties: (1) E_{∗,∗}^{∞} = {0}; (2) E_{0,j}^{2} ≅ \overline{H}_{j}(L; N);

(3) E_{i,0}^{1} ≅ H_{i}(A; N (A));

(4) For i, j > 0, dim_{G}(E_{i,j}^{1}) = 0.

(The fourth of these properties follows from Corollary 7.) It may be shown that for 1 < s ≤ r,

dim_{G}(E_{0,r−1}^{s}) = dim_{G}(E_{0,r−1}^{2}) and dim_{G}(E_{r,0}^{s}) = dim_{G}(E_{r,0}^{1}).

Since the E^{∞}-page is identically zero, the map d^{r} : E_{r,0}^{r} → E_{0,r−1}^{r} is an isomorphism, and hence dim_{G}(E_{r,0}^{r}) = dim_{G}(E_{0,r−1}^{r}). □

Remark. In the case when the Artin group A contains no finite conjugacy classes, the von Neumann dimension of a (finitely generated) Hilbert A-module determines the module up to isomorphism. Hence in this case, Theorem 21 is equivalent to Theorem 1.

10.2. ℓ^{p}-cohomology. For 1 ≤ p < ∞, let ℓ^{p}(G) denote the space of real-valued functions f on G for which ∑_{g∈G} |f(g)|^{p} is finite. Each ℓ^{p}(G) is a Banach space with respect to the obvious norm. If Y′ is a free G-CW-complex with only finitely many orbits of cells of each dimension, one may define reduced ℓ^{p}-cohomology groups H_{(p)}^{∗}(Y′) in the same way as the reduced ℓ^{2}-cohomology groups were defined in Section 3. For 1 < p < ∞, Gromov has proved vanishing theorems that imply the ℓ^{p}-analogues of Corollaries 7 and 8 in [16].

The main problem with attempting to generalize Theorems 1 and 3 to ℓ^{p}-cohomology is that there is no analogue of von Neumann dimension in this generality, and we do not know whether the analogue of Lemma 5 holds. Nevertheless, we conjecture that Theorems 1 and 3 remain valid for reduced ℓ^{p}-cohomology, for 1 < p < ∞.

References

[1] M. Bestvina and N. Brady, Morse theory and finiteness properties of groups, Invent. Math. 129 (1997), 445–470.

[2] N. Bourbaki, Groupes et Algèbres de Lie, Chapitres 4–6, Masson, Paris, 1981.

[3] M. Bridson and A. Haefliger, Metric Spaces of Non-positive Curvature, Springer-Verlag, Berlin and New York, 1999.

[4] E. Brieskorn, Sur les groupes de tresses, Sém. Bourbaki (1971/72), Springer Lec. Notes in Math. 317, Springer-Verlag, Berlin and New York, 1973, pp. 21–44.

[5] K. S. Brown, Cohomology of Groups, Springer-Verlag, Berlin and New York, 1982.

[6] R. Charney and M.W. Davis, The K(π, 1)-problem for hyperplane complements associated to infinite reflection groups, J. Amer. Math. Soc. 8 (1995), 597–627.

[7] R. Charney and M.W. Davis, Finite K(π, 1) s for Artin groups, Prospects in Topology (ed. F. Quinn), Annals of Math. Studies 138, Princeton Univ. Press, Princeton, 2000, pp. 110–124.

[8] J. Cheeger and M. Gromov, L_{2}-cohomology and group cohomology, Topology 25 (1986), 189–215.

[9] M.W. Davis, Nonpositive curvature and reflection groups, Handbook of Geometric Topology (eds. R. Daverman and R. Sher), Elsevier, Amsterdam, 2002, pp. 373-422.

[10] M.W. Davis and B. Okun, Vanishing theorems and conjectures for the ℓ^{2}- homology of right-angled Coxeter groups, Geometry and Topology 5 (2001), 7–74.

[11] C. De Concini and M. Salvetti, Cohomology of Artin groups, Math. Res. Lett. 3 (1996), 293–297.

[12] P. Deligne, Les immeubles des groupes de tresses généralisés, Invent. Math. 17 (1972), 273–302.

[13] W. Dicks and I. J. Leary, Presentations for subgroups of Artin groups, Proc. Amer. Math. Soc. 127 343–348 (1999).

[14] B. Eckmann, Introduction to ℓ_{2}-methods in topology: reduced ℓ_{2}-homology, harmonic chains, ℓ_{2}-Betti numbers, Israel J. Math. 117 (2000), 183–219

[15] M. Farber, Von Neumann Betti numbers and Novikov type inequalities, Proc. Amer. Math. Soc. 128 (2000), 2819–2827.

[16] M. Gromov, Asymptotic invariants of infinite groups, Geometric Group Theory 2 (eds. G. Niblo and M. Roller), London Math. Soc. Lecture Notes 182, Cambridge Univ. Press, Cambridge, 1993.

[17] A. Haefliger, Extensions of complexes of groups, Ann. Inst. Fourier (Grenoble) 42 (1992), 275–311.

[18] K. H. Kim and F. W. Roush, Homology of certain algebras defined by graphs, J. Pure Appl. Algebra 17, 179–186 (1980).

[19] W. Lück, L^{2}-invariants of regular coverings of compact manifolds and CWcomplexes, Handbook of Geometric Topology (eds. R. Daverman and R. Sher), Elsevier, Amsterdam, 2002, pp. 735–817.

[20] W. Lück, L^{2}-invariants and K-theory, Springer-Verlag, Berlin and New York, 2002.

[21] D. Quillen, Higher algebraic K-theory: I, Battelle Institute Conf., 1972, Springer Lec. Notes in Math. 341, Springer-Verlag, Berlin and New York, 1973, pp. 77–139.

[22] M. Salvetti, Topology of the complement of real hyperplanes in ℂ^{n}, Invent. Math. 88 (1987), 603–618.

[23] M. Salvetti, The homotopy type of Artin groups, Math. Res. Lett. 1 (1994), 565–577.

[24] C. A. Weibel, An Introduction to Homological Algebra, Cambridge Studies in Advanced Mathematics 38, Cambridge Univ. Press, Cambridge, 1994.

Department of Mathematics, The Ohio State University, 231 W. 18th Avenue, Columbus, Ohio 43210, United States

E-mail address: mdavis@math.ohio-state.edu

Faculty of Mathematical Studies, University of Southampton, Southampton, SO17 1BJ, United Kingdom

E-mail address: I.J.Leary@maths.soton.ac.uk